OpenStudy (anonymous):

If f(x) = 5x + 3, then what is f^–1(x) ?

OpenStudy (anonymous):

f(x)=5x+3 y=5x+3 Now interchange x and y x = 5y+ 3 5y= x-3 y= (x-3)/5